1-2 + 3 -4 + ....-98 + 99 equals
a) -50 b) -49 c) 0 d) 49 e) 50

Answer by Alan3354(21550)

1-2 + 3 -4 + ....-98 + 99 equals
a) -50 b) -49 c) 0 d) 49 e) 50
----------------
There are 99 numbers.
Put them in pairs,
1+99 = 100
-2-98 = -100
...
49+51 = 100 It's all even, at zero, to this point
Then the -50 has nothing to pair with
--> = -50